Doc. Math. J. DMV 643

Section 17 Control Theory and Optimization

In case of several authors, Invited Speakers are marked with a ^{∗}. David Applegate, Robert Bixby, Vašek Chv’atal and

William Cook^{∗}: On the Solution of Traveling Salesman Problems III 645 Michel X. Goemans: Semidefinite Programming and Combinatorial

Optimization . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . III 657 Richard H. Byrd and Jorge Nocedal^{∗}: Active Set and Interior

Methods for Nonlinear Optimization . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . III 667 Ranga Anbil, John J. Forrest and William R. Pulleyblank^{∗}:

Column Generation and the Airline Crew Pairing Problem . . . . . . . . . III 677 Alexander Schrijver: Routing and Timetabling by Topological

Search . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . III 687

Jan C. Willems: Open Dynamical Systems and their Control . . . . . . . III 697
Michal Kočvara and Jochem Zowe^{∗}: Free Material Optimization III 707

Documenta Mathematica · Extra Volume ICM 1998 · III ·

644

Documenta Mathematica · Extra Volume ICM 1998 · III ·